A metal pipe 90 cm long is cut into two pieces so that the longer piece is 15 cm longer than twice the length of the shorter piece. What is the length of the longer piece?

Answer by checkley79(3341)

X+(2X+15)=90
3X+15=90
3X=90-15
3X=75
X=75/3
X=25 cm. ANS FOR THE SHORTER PIECE.
2*25+15=50+15=65 ANS. FOR THE LONGER PIECE.
PROOF:
25+65=90
90=90